# Premier-TACO is a Few-Shot Policy Learner: Pretraining Multitask Representation via Temporal Action-Driven Contrastive Loss

### v3

@@ 1 Introduction
In the dynamic and ever-changing world we inhabit, the importance of sequential decision-making (SDM) in machine learning cannot be overstated. Unlike static tasks, sequential decisions reflect the fluidity of real-world scenarios, from robotic manipulations to evolving healthcare treatments. Just as foundation models in language, such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2019; Brown et al., 2020), have revolutionized natural language processing by leveraging vast amounts of textual data to understand linguistic nuances, pretrained foundation models hold similar promise for sequential decision-making (SDM). In language, these models capture the essence of syntax, semantics, and context, serving as a robust starting point for a myriad of downstream tasks. Analogously, in SDM, where decisions are influenced by a complex interplay of past actions, current states, and future possibilities, a pretrained foundation model can provide a rich, generalized understanding of decision sequences. This foundational knowledge, built upon diverse decision-making scenarios, can then be fine-tuned to specific tasks, much like how language models are adapted to specific linguistic tasks.
@@
Premier-TACO markedly enhances the effectiveness and efficiency of the temporal action contrastive learning (TACO) objective, as detailed in Zheng et al. (2023a), which delivers state-of-the-art outcomes in visual control tasks within a single-task setting. It extends these capabilities to efficient, large-scale multitask offline pretraining, broadening its applicability and performance. Specifically, while TACO considers every data point in a batch as a potential negative example, Premier-TACO strategically samples a single negative example from a proximate window of the subsequent state. This method ensures the negative example is visually akin to the positive one, necessitating that the latent representation captures control-relevant information, rather than relying on extraneous features like visual appearance. This efficient negative example sampling strategy adds no computational burden and is compatible with smaller batch sizes. In particular, on MetaWorld, using a batch size of for TACO, Premier-TACO achieves a 25% relative performance improvement. Premier-TACO can be seamlessly scaled for multitask offline pretraining, enhancing its usability and effectiveness.
@@
(2) [(W1) Versatility (W2) Efficiency] Through extensive empirical evaluation, we verify the effectiveness of Premier-TACO’s pretrained visual representations for few-shot learning on unseen tasks. On MetaWorld (Yu et al., 2019) and LIBERO Liu et al. (2023), with 5 expert trajectories, Premier-TACO outperforms the best baseline pretraining method by 37% and 17% respectively. Remarkably, in LIBERO, we are the first method to demonstrate benefits from pretraining. On Deepmind Control Suite (DMC) (Tassa et al., 2018), using only 20 trajectories, which is considerably fewer demonstrations than (Sun et al., 2023; Majumdar et al., 2023), Premier-TACO achieves the best performance across 10 challenging tasks, including the hard Dog and Humanoid tasks. This versatility extends even to unseen embodiments in DMC as well as unseen tasks with unseen camera views in MetaWorld.
@@ 2.2 TACO: Temporal Action Driven Contrastive Learning Objective
Temporal Action-driven Contrastive Learning (TACO) (Zheng et al., 2023a) is a reinforcement learning algorithm proposed for addressing the representation learning problem in visual continuous control. It aims to maximize the mutual information between representations of current states paired with action sequences and representations of the corresponding future states:
@@ 3 Method
We introduce Premier-TACO, a generalized pre-training approach specifically formulated to tackle the multi-task pre-training problem, enhancing sample efficiency and generalization ability for downstream tasks. Building upon the success of temporal contrastive loss, exemplified by TACO (Zheng et al., 2023a), in acquiring latent state representations that encapsulate individual task dynamics, our aim is to foster representation learning that effectively captures the intrinsic dynamics spanning a diverse set of tasks found in offline datasets. Our overarching objective is to ensure that these learned representations exhibit the versatility to generalize across unseen tasks that share the underlying dynamic structures.
Nevertheless, when adapted for multitask offline pre-training, the online learning objective of TACO (Zheng et al., 2023a) poses a notable challenge. Specifically, TACO’s mechanism, which utilizes the InfoNCE (van den Oord et al., 2019) loss, categorizes all subsequent states in the batch as negative examples. While this methodology has proven effective in single-task reinforcement learning scenarios, it encounters difficulties when extended to a multitask context. During multitask offline pretraining, image observations within a batch can come from different tasks with vastly different visual appearances, rendering the contrastive InfoNCE loss significantly less effective.
@@ 4 Experiment
Deepmind Control Suite (DMC): We consider a selection of 16 challenging tasks from Deepmind Control Suite. Note that compared with prior works such as Majumdar et al. (2023); Sun et al. (2023), we consider much harder tasks, including ones from the humanoid and dog domains, which feature intricate kinematics, skinning weights and collision geometry. For pretraining, we select six tasks (DMC-6), including Acrobot Swingup, Finger Turn Hard, Hopper Stand, Walker Run, Humanoid Walk, and Dog Stand. We generate an exploratory dataset for each task by sampling trajectories generated in exploratory stages of a DrQ-v2 (Yarats et al., 2022) learning agent. In particular, we sample 1000 trajectories from the online replay buffer of DrQ-v2 once it reaches the convergence performance. This ensures the diversity of the pretraining data, but in practice, such a high-quality dataset could be hard to obtain. So, later in the experiments, we will also relax this assumption and consider pretrained trajectories that are sampled from uniformly random actions. In terms of the encoder architecture, we pretrain Premier-TACO with the same shallow ConvNet encoder as in DrQv2 (Yarats et al., 2022).
@@
| Tasks |
LfS |
SMART |
Best PVRs |
TD3+BC |
Inverse |
CURL |
ATC |
SPR |
TACO |
Premier-TACO | |
| Seen Embodiments |
Finger Spin |
Hopper Hop |
Walker Walk |
Humanoid Walk |
Dog Trot |
| Unseen Embodiments |
Cup Catch |
Reacher Hard |
Cheetah Run |
Quadruped Walk |
Quadruped Run |
@@
| Unseen Tasks |
LfS |
SMART |
Best PVRs |
TD3+BC |
Inverse |
CURL |
ATC |
SPR |
TACO |
Premier-TACO |
Bin Picking |
Disassemble |
Hand Insert |
Peg Insert Side |
Pick Out Of Hole |
Pick Place Wall |
Shelf Place |
Stick Pull |
@@
Learn from Scratch: Behavior Cloning with randomly initialized shallow ConvNet encoder. We carefully implement the behavior cloning from scratch baseline. For DMC and MetaWorld, following Hansen et al. (2022a), we include the random shift data augmentation into behavior cloning. For LIBERO, we take the ResNet-T model architecture in Liu et al. (2023), which uses a transformer decoder module on top of the ResNet encoding to extract temporal information from a sequence of observations, addressing the non-Markovian characteristics inherent in human demonstration.
@@
Pretrained feature representation by Premier-TACO facilitates effective few-shot adaptation to unseen tasks. We measure the performance of pretrained visual representation for few-shot imitation learning of unseen downstream tasks in both DMC and MetaWorld. In particular, for DMC, we use 20 expert trajectories for imitation learning except for the two hardest tasks, Humanoid Walk and Dog Trot, for which we use 100 trajectories instead. Note that we only use of the number of expert trajectories used in Majumdar et al. (2023) and of those used in Sun et al. (2023).
@@
Ablation Study - Batch Size: Compared with TACO, the negative example sampling strategy employed in Premier-TACO allows us to sample harder negative examples within the same episode as the positive example. This implies the promising potential to significantly improve the performance of existing pretrained models across diverse domains. The full results of finetuning on all 18 tasks including Deepmind Control Suite and MetaWorld are in Appendix B.1. We expect Premier-TACO to work much better with small batch sizes, compared with TACO where the negative examples from a given batch could be coming from various tasks and thus the batch size required would scale up linearly with the number of pretraining tasks. In ours previous experimental results, Premier-TACO is pretrained with a batch size of 4096, a standard batch size used in contrastive learning literature. Here, to empirically verify the effects of different choices of the pretraining batch size, we train Premier-TACO and TACO with different batch sizes and compare their few-shot imitation learning performance. Figure 9 (left) displays the average performance of few-shot imitation learning across all ten tasks in the DeepMind Control Suite. As depicted in the figure, our model significantly outperform TACO across all batch sizes tested in the experiments, and exhibits performance saturation beyond a batch size of 4096. This observation substantiate that the negative example sampling strategy employed by Premier-TACO is indeed the key for the success of multitask offline pretraining.
@@ 5 Related Work
Existing works, including R3M (Nair et al., 2022), VIP (Ma et al., 2023), MVP (Xiao et al., 2022), PIE-G (Yuan et al., 2022), and VC-1 (Majumdar et al., 2023), focus on self-supervised pre-training for building foundation models but struggle with the domain gap in sequential decision-making tasks. Recent studies, such as one by Hansen et al. (2022a), indicate that models trained from scratch often outperform pre-trained representations. Approaches like SMART (Sun et al., 2023) and DualMind (Wei et al., 2023) offer control-centric pre-training, but at the cost of extensive fine-tuning or task sets. Contrastive learning techniques like CURL (Laskin et al., 2020), CPC (Henaff, 2020), ST-DIM (Anand et al., 2019), and ATC (Stooke et al., 2021a) have succeeded in visual RL, but mainly focus on high-level features and temporal dynamics without a holistic consideration of state-action interactions, a gap partially filled by TACO (Zheng et al., 2023a). Our work builds upon these efforts but eliminates the need for extensive task sets and fine-tuning, efficiently capturing control-relevant features. This positions our method as a distinct advancement over DRIML (Mazoure et al., 2020) and Homer (Misra et al., 2019), which require more computational or empirical resources.
@@ 6 Conclusion
7 Acknowledgement
This work is partially done during Ruijie Zheng’s internship at Microsoft Research. Zheng, Wang, and Huang are supported by National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships.
@@ Appendix A Detailed Discussion of Related Work
For control tasks, several pretraining frameworks have emerged to model state-action interactions from high-dimensional observations by leveraging causal attention mechanisms. SMART (Sun et al., 2023) introduces a self-supervised and control-centric objective to train transformer-based models for multitask decision-making, although it requires additional fine-tuning with large number of demonstrations during downstream time. As an improvement, DualMind (Wei et al., 2023) pretrains representations using 45 tasks for general-purpose decision-making without task-specific fine-tuning. Besides, some methods (Sekar et al., 2020; Mendonca et al., 2021; Yarats et al., 2021; Sun et al., 2022) first learn a general representation by exploring the environment online, and then use this representation to train the policy on downstream tasks. In comparison, our approach is notably more efficient and doesn’t require training with such an extensive task set. Nevertheless, we provide empirical evidence demonstrating that our method can effectively handle multi-task pretraining.
@@
Contrastive Representation for Visual RL Visual RL (Hafner et al., 2020; Yarats et al., 2022; Hansen et al., 2022c; Zheng et al., 2023b; Wang et al., 2024; Xu et al., 2024) is long-standing challenge due to the entangled problem of representation learning and credit assignment. Contrastive learning plays a pivotal role in enhancing sample efficiency of visual RL by leveraging similarity constraints between data to learn effective visual representations. CURL (Laskin et al., 2020) extracts high-level features by utilizing InfoNCE(van den Oord et al., 2019) to maximize agreement between augmented observations, although it does not explicitly consider temporal relationships between states. Several approaches, such as CPC (Henaff, 2020), ST-DIM (Anand et al., 2019), and ATC (Stooke et al., 2021a) , introduce temporal dynamics into the contrastive loss. They do so by maximizing mutual information between states with short temporal intervals, facilitating the capture of temporal dependencies. DRIML (Mazoure et al., 2020) proposes a policy-dependent auxiliary objective that enhances agreement between representations of consecutive states, specifically considering the first action of the action sequence. Recent advancements by Kim et al. (2022); Zhang et al. (2021) incorporate actions into the contrastive loss, emphasizing behavioral similarity. TACO (Zheng et al., 2023a) takes a step further by learning both state and action representations. It optimizes the mutual information between the representations of current states paired with action sequences and the representations of corresponding future states. In our approach, we build upon the efficient extension of TACO, harnessing the full potential of state and action representations for downstream tasks. On the theory side, the Homer algorithm (Misra et al., 2019) uses a binary temporal contrastive objective reminiscent of the approach used here, which differs by abstracting actions as well states, using an ancillary embedding, removing leveling from the construction, and of course extensive empirical validation.
@@
Hard Negative Sampling Strategy in Contrastive Learning Our proposed negative example sampling strategy in Premier-TACO is closely related to hard negative example mining in the literature of self-supervised learning as well as other areas of machine learning. Hard negative mining is indeed used in a variety of tasks, such as facial recognition (Wan et al., 2016), object detection (Shrivastava et al., 2016), tracking (Nam & Han, 2016), and image-text retrieval (Pang et al., 2019; Li et al., 2021), by introducing negative examples that are more difficult than randomly chosen ones to improve the performance of models. Within the regime of self-supervised learning, different negative example sampling strategies have also been discussed both empirically and theoretically to improve the quality of pretrained representation. In particular, Robinson et al. (2021) modifies the original NCE objective by developing a distribution over negative examples, which prioritizes pairs with currently similar representations. Kalantidis et al. (2020) suggests to mix hard negative examples within the latent space. Ma et al. (2021) introduce a method to actively sample uncertain negatives by calculating the gradients of the loss function relative to the model’s most confident predictions. Furthermore, Tabassum et al. (2022) samples negatives that combine the objectives of identifying model-uncertain negatives, selecting negatives close to the anchor point in the latent embedding space, and ensuring representativeness within the sample population.
While our proposed approach bears some resemblance to existing negative sampling strategies in contrastive learning literature, we are dealing with unique challenges in sequential decision making, such as data distribution shift, task heterogeneity, and data qualities, as discussed in the introduction. Advancing the work of TACO (Zheng et al., 2023a), which is specifically designed to capture the control-relevant information in latent representation, Premier-TACO introduces a straightforward yet effective negative sampling strategy. Tailored toward multitask representation pretraining, this strategy involves sampling one negative example from a window centered around the anchor point, which is both computationally efficient and demonstrates superior performance in few-shot adaptation.
Appendix B Additional Experiments Results

### v4

@@ 1 Introduction
In the dynamic and ever-changing world we inhabit, the importance of sequential decision-making (SDM) in machine learning cannot be overstated. Unlike static tasks, sequential decisions reflect the fluidity of real-world scenarios, from robotic manipulations to evolving healthcare treatments. Just as foundation models in language, such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2019; Brown et al., 2020), have revolutionized natural language processing by leveraging vast amounts of textual data to understand linguistic nuances, pretrained foundation models hold similar promise for sequential decision-making (SDM). In language, these models capture the essence of syntax, semantics, and context, serving as a robust starting point for a myriad of downstream tasks. Analogously, in SDM, where decisions are influenced by a complex interplay
of past actions, current states, and future possibilities, a pretrained foundation model can provide a rich, generalized understanding of decision sequences. This foundational knowledge, built upon diverse decision-making scenarios, can then be fine-tuned to specific tasks, much like how language models are adapted to specific linguistic tasks.
@@
Premier-TACO markedly enhances the effectiveness and efficiency of the temporal action contrastive learning (TACO) objective, as detailed in Zheng et al. (2023), which delivers state-of-the-art outcomes in visual control tasks within a single-task setting. It extends these capabilities to efficient, large-scale multitask offline pretraining, broadening its applicability and performance. Specifically, while TACO considers every data point in a batch as a potential negative example, Premier-TACO strategically samples a single negative example from a proximate window of the subsequent state. This method ensures the negative example is visually akin to the positive one, necessitating that the latent representation captures control-relevant information, rather than relying on extraneous features like visual appearance. This efficient negative example sampling strategy adds no computational burden and is compatible with smaller batch sizes. In particular, on MetaWorld, using a batch size of for TACO, Premier-TACO achieves a 25% relative performance improvement. Premier-TACO can be seamlessly scaled for multitask offline pretraining, enhancing its usability and effectiveness.
@@
(2) [(W1) Versatility (W2) Efficiency] Through extensive empirical evaluation, we verify the effectiveness of Premier-TACO’s pretrained visual representations for few-shot learning on unseen tasks. On MetaWorld (Yu et al., 2019) and LIBERO (Liu et al., 2023), with 5 expert trajectories, Premier-TACO outperforms the best baseline pretraining method by 37% and 17% respectively. Remarkably, in LIBERO, we are the first method to demonstrate benefits from pretraining. On Deepmind Control Suite (DMC) (Tassa et al., 2018), using only 20 trajectories, which is considerably fewer demonstrations than (Sun et al., 2023; Majumdar et al., 2023), Premier-TACO achieves the best performance across 10 challenging tasks, including the hard Dog and Humanoid tasks. This versatility extends even to unseen embodiments in DMC as well as unseen tasks with unseen camera views in MetaWorld.
@@ 2.2 TACO: Temporal Action Driven Contrastive Learning Objective
Temporal Action-driven Contrastive Learning (TACO) (Zheng et al., 2023) is a reinforcement learning algorithm proposed for addressing the representation learning problem in visual continuous control. It aims to maximize the mutual information between representations of current states paired with action sequences and representations of the corresponding future states:
@@ 3 Method
We introduce Premier-TACO, a generalized pre-training approach specifically formulated to tackle the multi-task pre-training problem, enhancing sample efficiency and generalization ability for downstream tasks. Building upon the success of temporal contrastive loss, exemplified by TACO (Zheng et al., 2023), in acquiring latent state representations that encapsulate individual task dynamics, our aim is to foster representation learning that effectively captures the intrinsic dynamics spanning a diverse set of tasks found in offline datasets. Our overarching objective is to ensure that these learned representations exhibit the versatility to generalize across unseen tasks that share the underlying dynamic structures.
Nevertheless, when adapted for multitask offline pre-training, the online learning objective of TACO (Zheng et al., 2023) poses a notable challenge. Specifically, TACO’s mechanism, which utilizes the InfoNCE (van den Oord et al., 2019) loss, categorizes all subsequent states in the batch as negative examples. While this methodology has proven effective in single-task reinforcement learning scenarios, it encounters difficulties when extended to a multitask context. During multitask offline pretraining, image observations within a batch can come from different tasks with vastly different visual appearances, rendering the contrastive InfoNCE loss significantly less effective.
@@ 4 Experiment
Deepmind Control Suite (DMC): We consider a selection of 16 challenging tasks from Deepmind Control Suite. Note that compared with prior works such as (Majumdar et al., 2023; Sun et al., 2023), we consider much harder tasks, including ones from the humanoid and dog domains, which feature intricate kinematics, skinning weights and collision geometry. For pretraining, we select six tasks (DMC-6), including Acrobot Swingup, Finger Turn Hard, Hopper Stand, Walker Run, Humanoid Walk, and Dog Stand. We generate an exploratory dataset for each task by sampling trajectories generated in exploratory stages of a DrQ-v2 (Yarats et al., 2022) learning agent. In particular, we sample 1000 trajectories from the online replay buffer of DrQ-v2 once it reaches the convergence performance. This ensures the diversity of the pretraining data, but in practice, such a high-quality dataset could be hard to obtain. So, later in the experiments, we will also relax this assumption and consider pretrained trajectories that are sampled from uniformly random actions. In terms of the encoder architecture, we pretrain Premier-TACO with the same shallow ConvNet encoder as in DrQv2 (Yarats et al., 2022).
@@
| Tasks | LfS | SMART | Best PVRs | TD3+BC | Inverse | CURL | ATC | SPR | TACO | Premier-TACO | |
| Seen Embodiments | Finger Spin | ||||||||||
| Hopper Hop | |||||||||||
| Walker Walk | |||||||||||
| Humanoid Walk | |||||||||||
| Dog Trot | |||||||||||
| Unseen Embodiments | Cup Catch | ||||||||||
| Reacher Hard | |||||||||||
| Cheetah Run | |||||||||||
| Quadruped Walk | |||||||||||
| Quadruped Run | |||||||||||
@@
| Unseen Tasks | LfS | SMART | Best PVRs | TD3+BC | Inverse | CURL | ATC | SPR | TACO | Premier-TACO |
| Bin Picking | ||||||||||
| Disassemble | ||||||||||
| Hand Insert | ||||||||||
| Peg Insert Side | ||||||||||
| Pick Out Of Hole | ||||||||||
| Pick Place Wall | ||||||||||
| Shelf Place | ||||||||||
| Stick Pull | ||||||||||
@@
Learn from Scratch: Behavior Cloning with randomly initialized shallow ConvNet encoder. We carefully implement the behavior cloning from scratch baseline. For DMC and MetaWorld, following (Hansen et al., 2022a), we include the random shift data augmentation into behavior cloning. For LIBERO, we take the ResNet-T model architecture in (Liu et al., 2023), which uses a transformer decoder module on top of the ResNet encoding to extract temporal information from a sequence of observations, addressing the non-Markovian characteristics inherent in human demonstration.
@@
Pretrained feature representation by Premier-TACO facilitates effective few-shot adaptation to unseen tasks. We measure the performance of pretrained visual representation for few-shot imitation learning of unseen downstream tasks in both DMC and MetaWorld. In particular, for DMC, we use 20 expert trajectories for imitation learning except for the two hardest tasks, Humanoid Walk and Dog Trot, for which we use 100 trajectories instead. Note that we only use of the number of expert trajectories used in (Majumdar et al., 2023) and of those used in (Sun et al., 2023).
@@
Ablation Study - Batch Size: Compared with TACO, the negative example sampling strategy employed in Premier-TACO allows us to sample harder negative examples within the same episode as the positive example. This implies the promising potential to significantly improve the performance of existing pretrained models across diverse domains. The full results of finetuning on all 18 tasks including Deepmind Control Suite and MetaWorld are in Appendix B.1. We expect Premier-TACO to work much better with small batch sizes, compared with TACO where the negative examples from a given batch could be coming from various tasks and thus the batch size required would scale up linearly with the number of pretraining tasks. In ours previous experimental results, Premier-TACO is pretrained with a batch size of 4096, a standard batch size used in contrastive learning literature. Here, to empirically verify the effects of different choices of the pretraining batch size, we train Premier-TACO and TACO with different batch sizes and compare their few-shot imitation learning performance.
Figure 9 (left) displays the average performance of few-shot imitation learning across all ten tasks in the DeepMind Control Suite. As depicted in the figure, our model significantly outperform TACO across all batch sizes tested in the experiments, and exhibits performance saturation beyond a batch size of 4096. This observation substantiate that the negative example sampling strategy employed by Premier-TACO is indeed the key for the success of multitask offline pretraining.
@@ 5 Related Work
Existing works, including R3M (Nair et al., 2022), VIP (Ma et al., 2023), MVP (Xiao et al., 2022), PIE-G (Yuan et al., 2022), and VC-1 (Majumdar et al., 2023), focus on self-supervised pre-training for building foundation models but struggle with the domain gap in sequential decision-making tasks. Recent studies, such as one by Hansen et al. (2022a), indicate that models trained from scratch often outperform pre-trained representations. Approaches like SMART (Sun et al., 2023) and DualMind (Wei et al., 2023) offer control-centric pre-training, but at the cost of extensive fine-tuning or task sets. Contrastive learning techniques like CURL (Laskin et al., 2020), CPC (Henaff, 2020), ST-DIM (Anand et al., 2019), and ATC (Stooke et al., 2021a) have succeeded in visual RL, but mainly focus on high-level features and temporal dynamics without a holistic consideration of state-action interactions, a gap partially filled by TACO (Zheng et al., 2023). Our work builds upon these efforts but eliminates the need for extensive task sets and fine-tuning, efficiently capturing control-relevant features. This positions our method as a distinct advancement over DRIML (Mazoure et al., 2020) and Homer (Misra et al., 2019), which require more computational or empirical resources.
@@ 6 Conclusion
Acknowledgements
Zheng, Wang, and Huang are supported by National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Detailed Discussion of Related Work
For control tasks, several pretraining frameworks have emerged to model state-action interactions from high-dimensional observations by leveraging causal attention mechanisms. SMART (Sun et al., 2023) introduces a self-supervised and control-centric objective to train transformer-based models for multitask decision-making, although it requires additional fine-tuning with large number of demonstrations during downstream time. As an improvement, DualMind (Wei et al., 2023) pretrains representations using 45 tasks for general-purpose decision-making without task-specific fine-tuning. Besides, some methods (Sekar et al., 2020; Mendonca et al., 2021; Yarats et al., 2021a; Sun et al., 2022) first learn a general representation by exploring the environment online, and then use this representation to train the policy on downstream tasks. In comparison, our approach is notably more efficient and doesn’t require training with such an extensive task set. Nevertheless, we provide empirical evidence demonstrating that our method can effectively handle multi-task pretraining.
@@
Contrastive Representation for Visual RL Visual RL (Yarats et al., 2021b, 2022; Hafner et al., 2020; Hansen et al., 2022b; Xu et al., 2024; Ji et al., 2024) is long-standing challenge due to the entangled problem of representation learning and credit assignment. In the context of visual reinforcement learning (RL), contrastive learning plays a pivotal role in training robust state representations from raw visual inputs, thereby enhancing sample efficiency. CURL (Laskin et al., 2020) extracts high-level features by utilizing InfoNCE(van den Oord et al., 2019) to maximize agreement between augmented observations, although it does not explicitly consider temporal relationships between states. Several approaches, such as CPC (Henaff, 2020), ST-DIM (Anand et al., 2019), and ATC (Stooke et al., 2021a) , introduce temporal dynamics into the contrastive loss. They do so by maximizing mutual information between states with short temporal intervals, facilitating the capture of temporal dependencies. DRIML (Mazoure et al., 2020) proposes a policy-dependent auxiliary objective that enhances agreement between representations of consecutive states, specifically considering the first action of the action sequence. Recent advancements by Kim et al. (2022); Zhang et al. (2021) incorporate actions into the contrastive loss, emphasizing behavioral similarity. TACO (Zheng et al., 2023) takes a step further by learning both state and action representations. It optimizes the mutual information between the representations of current states paired with action sequences and the representations of corresponding future states. In our approach, we build upon the efficient extension of TACO, harnessing the full potential of state and action representations for downstream tasks. On the theory side, the Homer algorithm (Misra et al., 2019) uses a binary temporal contrastive objective reminiscent of the approach used here, which differs by abstracting actions as well states, using an ancillary embedding, removing leveling from the construction, and of course extensive empirical validation.
@@
Hard Negative Sampling Strategy in Contrastive Learning Our proposed negative example sampling strategy in Premier-TACO is closely related to hard negative example mining in the literature of self-supervised learning as well as other areas of machine learning. Hard negative mining is indeed used in a variety of tasks, such as facial recognition (Wan et al., 2016), object detection (Shrivastava et al., 2016), tracking (Nam & Han, 2016), and image-text retrieval (Pang et al., 2019; Li et al., 2021), by introducing negative examples that are more difficult than randomly chosen ones to improve the performance of models. Within the regime of self-supervised learning, different negative example sampling strategies have also been discussed both empirically and theoretically to improve the quality of pretrained representation. In particular, (Robinson et al., 2021) modifies the original NCE objective by developing a distribution over negative examples, which prioritizes pairs with currently similar representations. (Kalantidis et al., 2020) suggests to mix hard negative examples within the latent space. (Ma et al., 2021) introduce a method to actively sample uncertain negatives by calculating the gradients of the loss function relative to the model’s most confident predictions. Furthermore, (Tabassum et al., 2022) samples negatives that combine the objectives of identifying model-uncertain negatives, selecting negatives close to the anchor point in the latent embedding space, and ensuring representativeness within the sample population.
Comparison with Offline Meta-RL Methods Compared with offline meta-rl methods, feature representation learning with self-supervised/contrastive objectives, such as Premier-TACO, can efficiently leverage low-quality datasets (e.g., datasets collected by rolling out random actions in the DeepMind Control Suite) (Xu et al., 2023; Gao et al., 2023; Mitchell et al., 2021). Offline meta RL, in contrast, relies on datasets with good coverage to learn effective policies and typically addresses tasks with smaller shifts between meta-training and meta-testing (e.g., varying velocities in MuJoCo’s halfcheetah).
In particular, HDT (Xu et al., 2023) utilizes a hyper-network as an adaptation module to encode expert demonstrations and augment the base Decision Transformer (Chen et al., 2021) model. Unlike HDT, Premier-TACO can adapt to unseen embodiments with different action spaces by initializing a new policy head. In contrast, HDT’s hyper-network architecture does not easily accommodate unseen action spaces without significant modifications. CSRO (Gao et al., 2023) focuses on smaller task distribution shift as in prior offline meta RL works, such as humanoid in MuJoCo running in different directions. The task representation learned in CSRO is also not able to handle the unseen action spaces. On the contrary, Premier-TACO tackles a broader problem and experimental setting, enabling our representation to generalize to unseen downstream tasks and unseen embodiments with new action spaces.
Other pretraining schemes for decision-making. In this work, we primarily focuses on pretraining visual state representations. Several other existing works aim to solve multitask pretraining for sequential decision making from a different perspective, through the discovery of temporal action abstractions (i.e. skills or options). These works propose to pretrain temporally extended action primitives and subsequently use them for shortening the effective decision-making horizon during high-level policy induction, including CompILE (Kipf et al., 2019), RPL (Gupta et al., 2019), OPAL (Ajay et al., 2021), LOVE (Jiang et al., 2023), and PRISE (Zheng et al., 2024). They often operate in two stages: learning the primitives during the first and applying them to solve a downstream task during the second, possibly adapting the primitives in the process. Compared with the visual state representation proposed in Premier-TACO, temporal action abstractions go in an orthogonal direction, and combining the benefits of both pretrained state representation and temporal action abstractions could be an exciting future direction.
Appendix B Additional Experiment Results